# Parameter-Efficient Fine-Tuning of State Space Models

### v1

@@ Acknowledgement
Appendix A In-depth Introduction of Baselines
@@
LoRA (Low-Rank Adaptation) focuses on fine-tuning large models by freezing most of the pretrained parameters and injecting trainable low-rank matrices into each layer of the Transformer’s architecture. The intuition behind using low-rank matrices comes from linear algebra, where a large matrix can be closely approximated by the product of two smaller matrices. The number of trainable parameters can be controlled with the rank of the low-rank matrices. LoRA also uses a scaling parameter (LoRA alpha) for the weight matrices to control the balance of the original model weights and LoRA weights during training. After fine-tuning, LoRA weights can be merged with the original model weights, introducing no additional inference overhead.
@@
Prefix-tuning also prepends trainable tokens to the input like prompt tuning but injects separate prefixes in every layer. For each Transformer layer, prefix-tuning prepends trainable embeddings to the attention’s ${\bm{K}}$ and ${\bm{V}}$ matrix. The authors have found that directly training these prefixes can lead to unstable training, so they propose to over-parameterize them with a large MLP to increase training stability. After training, the MLP can be dropped. Like prompt tuning, prefix-tuning introduces an inference overhead, scaling linearly with the number of trainable embeddings.
@@ Appendix B Details of Datasets
In this paper, we consider five datasets across three domains: (i) Natural Language Understanding (NLU), represented by GLUE (Wang et al., 2019); (ii) Natural Language Generation (NLG), including SAMSum (Gliwa et al., 2019), Spider (Yu et al., 2018) and DART (Nan et al., 2021); and (iii) Computer Vision (CV), represented by CIFAR-10 (Krizhevsky et al., 2009).
@@
The CIFAR-10 (Canadian Institute For Advanced Research) dataset is a collection of images that are commonly used to train machine learning and computer vision algorithms. It is one of the most widely used datasets for image classification. The CIFAR-10 dataset contains 60,000 (50,000 for training, 10,000 for validation) 32$\times$32 color images in 10 different classes. The 10 different classes are: airplane, car, bird, cat, deer, dog, frog, horse, ship, and truck. There are 6,000 images of each class. For training, we center crop each image to 24$\times$24 pixels and flatten each image to a string, with a total of 24$\times$24$\times$3 words, where each word is a number between 0-255 representing the respective pixel value. Although CIFAR-10 is a dataset for computer vision, previous work (Dinh et al., 2022) showed that Transformers can be adapted to the vision domain from the language domain, and we tested this ability on the state-space model.
Data Size (Train) Size (Val) Size (Test) Max. seq. len. #Epochs Mamba Size Metrics
GLUE RTE 1992 498 277 291 10 130M Accuracy
MRPC 2934 734 408 105 10 130M Accuracy
CoLA 6840 1711 1043 47 10 130M Matthews corr.
SST-2 53879 13470 872 68 10 130M Accuracy
QNLI 83794 20949 5463 602 10 130M Accuracy
QQP 291076 72770 40430 316 3 130M Accuracy
MNLI 314161 78541 19647 425 3 130M Accuracy
Spider 5543 1375 1034 1412 10 1.4B, 2.8B Accuracy
SAMSum 14732 818 819 1174 10 1.4B ROUGE
DART 62659 2768 5097 491 10 130M METEOR, BLEU
CIFAR-10 40000 10000 10000 1730 5 130M Accuracy
@@ C.2 Extended Results on Benchmarking Existing PEFT Methods
We present comprehensive fine-tuning results for the GLUE benchmark (Wang et al., 2019), DART dataset (Nan et al., 2021), SAMSum dataset (Gliwa et al., 2019) and Spider dataset (Yu et al., 2018) in Table 7, Table 8, Table 9 and Table 10, respectively. These experimental results encompass various LoRA implementations (on different weight matrices and modules) and provide more fine-grained results across all subtasks.
@@
Table 7: Full experimental results on the GLUE (Wang et al., 2019) benchmark. We report accuracy ($\uparrow$) for RTE, MRPC, SST-2, QNLI, QQP, and MNLI tasks. CoLA performance is measured using Matthews Correlation Coefficient ($\uparrow$). Mamba-130M is employed in this experiment. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. ${\bm{\beta}}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
@@
Table 8: Full experimental results on the DART (Nan et al., 2021) benchmark. We report METEOR ()($\uparrow$)( ↑ ) and BLEU ()($\uparrow$)( ↑ ) scores. Mamba-130M is utilized in this experiment. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. ${\bm{\beta}}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
@@
Table 9: Full experimental results on the SAMSum (Gliwa et al., 2019) benchmark. R1, R2, and RL represent ROUGE-1 ($\uparrow$), ROUGE-2 ($\uparrow$), and ROUGE-L ($\uparrow$), respectively. Mamba-1.4B is utilized in this experiment. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. ${\bm{\beta}}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
@@
(a) Comprehensive experimental results on Spider using Mamba-1.4B.
@@
(b) Comprehensive experimental results on Spider using Mamba-2.8B.
Table 10: Full experimental results on Spider (Yu et al., 2018) dataset. We report the accuracy ($\uparrow$) for Spider and its subsets. We consider two models in our experiments: Mamba-1.4B and Mamba-2.8B. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. ${\bm{\beta}}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
@@
Table 11: Full experimenal results on the CIFAR-10 (Krizhevsky et al., 2009) dataset. We report accuracy ($\uparrow$). Mama-130M is utilized in this experiment. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. ${\bm{\beta}}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
C.3 Limitations of Applying Prompt-based Methods on SSMs

### v3

@@ Acknowledgment
Appendix A Additional Related Works
A.1 Additional Related Works on SSMs
Despite of the remarkable performance of SSMs on certain tasks of sequence modeling, SSMs still showed worse performance than Transformers on language modeling. Fu et al. (2022) transitioned from synthetic language modeling tasks to real language modeling tasks with SSMs. They proposed H3, which is inspired by Linear Attention (Katharopoulos et al., 2020), introducing both diagonal SSM and shift SSM. Recently, Mamba (Gu & Dao, 2024; Dao & Gu, 2024) escaped from linear time invariance (LTI) modeling by introducing input-dependent terms and achieved better performance than Transformers on language modeling. Furthermore, several hybrid models (Lieber et al., 2025; Park et al., 2024) tried to exploit the advantages of both SSMs and Transformers.
A.2 Additional Related Works on PEFT
@@
LoRA (Low-Rank Adaptation) focuses on fine-tuning large models by freezing pretrained parameters and injecting trainable low-rank matrices into each layer of the Transformer architecture. The intuition behind using low-rank matrices comes from linear algebra, where a large matrix can be closely approximated by the product of two smaller matrices. The number of trainable parameters can be controlled with the rank of the low-rank matrices. LoRA also uses a scaling parameter (LoRA alpha) for the weight matrices to control the balance of the original model weights and LoRA weights during training. After fine-tuning, LoRA weights can be merged with the original model weights, introducing no additional inference overhead.
@@
Prefix-tuning also prepends trainable tokens to the input like prompt tuning but injects separate prefixes in every layer. For each Transformer layer, prefix-tuning prepends trainable embeddings to the attention’s ${\bm{K}}$ and ${\bm{V}}$ matrix. The authors have found that directly training these prefixes can lead to unstable training, so they propose to over-parameterize them with a large MLP to increase training stability. After training, the MLP can be dropped. Like prompt tuning, prefix-tuning introduces an inference overhead, scaling with the number of trainable embeddings.
@@
Theoretical Understanding of PEFT.
Numerous efforts have been made to theoretically understand existing PEFT methods. For input-injection methods, Wang et al. (2023b), Petrov et al. (2024), and Oymak et al. (2023) have theoretically analyzed the effectiveness and limitations of prompt tuning and prefix-tuning for Transformer-based models. For LoRA, Zeng & Lee (2024) explored its expressive power by demonstrating that even a randomly initialized model can be adapted to match any smaller target model using LoRA. Some of our theoretical analysis draws upon the framework established by Zeng & Lee (2024). Jang et al. (2024) conducted a theoretical exploration of LoRA within the neural tangent kernel (NTK) regime.
@@ Appendix B Details of Datasets
In this paper, we consider six datasets across three domains: (i) Natural Language Understanding (NLU), represented by GLUE (Wang et al., 2019); (ii) Natural Language Generation (NLG), including SAMSum (Gliwa et al., 2019), Spider (Yu et al., 2018) and DART (Nan et al., 2021); and (iii) Computer Vision (CV), represented by CIFAR-10 (Krizhevsky et al., 2009) and CelebA (Liu et al., 2015).
@@
The CIFAR-10 (Canadian Institute For Advanced Research) dataset is a collection of images that are commonly used to train machine learning and computer vision algorithms. It is one of the most widely used datasets for image classification. The CIFAR-10 dataset contains 60,000 (50,000 for training, 10,000 for validation) 32$\times$32 color images in 10 different classes. The 10 different classes are: airplane, car, bird, cat, deer, dog, frog, horse, ship, and truck. There are 6,000 images of each class. For training, we center crop each image to 24$\times$24 pixels and flatten each image to a string, with a total of 24$\times$24$\times$3 words, where each word is a number between 0-255 representing the respective pixel value. Although CIFAR-10 is a dataset for computer vision, previous work (Dinh et al., 2022) showed that Transformers can be adapted to the vision domain from the language domain. In our work, we extend this investigation to SSMs, examining their ability to perform on vision data.
CelebA (Liu et al., 2015).
The CelebA (CelebFaces Attributes) dataset is an extensive collection of more than 200,000 celebrity images, each tagged with 40 attributes. This dataset is notable for its diversity, volume, and comprehensive annotations, encompassing 10,177 distinct identities, 202,599 facial images, and annotations of five landmark points with 40 binary attributes per image. The dataset, which includes images with varied poses and complex backgrounds, is an essential resource for tasks in computer vision such as face recognition, attribute analysis, and detection, as well as facial landmark localization, and it offers significant utility in face editing and synthesis.
Dataset Size (Train) Size (Val) Size (Test) Max. seq. len. #Epochs Mamba Size Jamba Size Metrics
GLUE RTE 1992 498 277 291 10 130M 319M Accuracy
MRPC 2934 734 408 105 10 130M 319M Accuracy
CoLA 6840 1711 1043 47 10 130M 319M Matthews corr.
SST-2 53879 13470 872 68 10 130M 319M Accuracy
QNLI 83794 20949 5463 602 10 130M 319M Accuracy
QQP 291076 72770 40430 316 3 130M 319M Accuracy
MNLI 314161 78541 19647 425 3 130M 319M Accuracy
Spider 5543 1375 1034 1412 10 1.4B, 2.8B 52B Accuracy
SAMSum 14732 818 819 1174 10 1.4B 52B ROUGE
DART 62659 2768 5097 491 10 130M 52B METEOR, BLEU
CIFAR-10 40000 10000 10000 1730 5 130M 319M Accuracy
CelebA 162770 19867 19962 12614 3 130M 319M Accuracy
@@ C.2 Extended Results on Benchmarking Existing PEFT Methods
Mamba-I.
We present comprehensive fine-tuning results for the GLUE benchmark (Wang et al., 2019), DART dataset (Nan et al., 2021), SAMSum dataset (Gliwa et al., 2019), Spider dataset (Yu et al., 2018), and CIFAR-10 (Krizhevsky et al., 2009) in Table 6, Table 7, Table 8, Table 9, and Table 10 respectively. These experimental results encompass various LoRA implementations (on different weight matrices and modules) and provide more fine-grained results across all subtasks.
Mamba-II.
Table 11 and Table 12 present the benchmark results of LoRA and full fine-tuning across different layers of Mamba-II. We follow the same experimental setup used for Mamba-I and demonstrate that, on Mamba-II, our conclusion holds: LoRA is more effective on linear projection layers than on SSM modules.
Jamba.
Table 13 presents the benchmark results of LoRA and full fine-tuning across different layers of Jamba. Our findings demonstrate that, on Jamba, LoRA is more effective on linear projection layers than on SSM modules, which aligns with our conclusion on Mamba.
@@
Table 6: Full benchmark results on the GLUE (Wang et al., 2019) benchmark using Mamba-I 130M. We report accuracy ($\uparrow$) for RTE, MRPC, SST-2, QNLI, QQP, and MNLI tasks. CoLA performance is measured using Matthews Correlation Coefficient ($\uparrow$). In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for the SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. $\bm{\beta}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
@@
Table 7: Full benchmark results on the DART (Nan et al., 2021) benchmark using Mamba-I 130M. We report METEOR ()($\uparrow$)( ↑ ) and BLEU ()($\uparrow$)( ↑ ) scores. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. $\bm{\beta}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
@@
Table 8: Full benchmark results on the SAMSum (Gliwa et al., 2019) benchmark using Mamba-I 1.4B. R1, R2, and RL represent ROUGE-1 ($\uparrow$), ROUGE-2 ($\uparrow$), and ROUGE-L ($\uparrow$), respectively. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. $\bm{\beta}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
@@
(a) Full benchmark results on Spider using Mamba-I 1.4B.
@@
(b) Full benchmark results on Spider using Mamba-I 2.8B.
Table 9: Full benchmark results on Spider (Yu et al., 2018) dataset using Mamba-I. We report the accuracy ($\uparrow$) for Spider and its subsets. We consider two models in our experiments: Mamba-I 1.4B and Mamba-I 2.8B. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. $\bm{\beta}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
@@
Table 10: Full benchmark results on the CIFAR-10 (Krizhevsky et al., 2009) dataset using Mamba-I 130M. We report accuracy ($\uparrow$). In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. $\bm{\beta}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
Layer Method # Params (%) METEOR BLEU
All All Full $100.0$ 66.565 710 306 167 666.565710306167666.565\,710\,306\,167\,666.565 710 306 167 6 34.870 043 396 949 7734.8700433969497734.870\,043\,396\,949\,7734.870 043 396 949 77
LoRA 1.393 771 181 403 076 71.39377118140307671.393\,771\,181\,403\,076\,71.393 771 181 403 076 7 66.923 677 921 295 1766.9236779212951766.923\,677\,921\,295\,1766.923 677 921 295 17 45.405 793 190 002 4445.4057931900024445.405\,793\,190\,002\,4445.405 793 190 002 44
Linear Projection Matrices $\bm{W}_{\text{in}},\bm{W}_{\text{out}}$ LoRA 1.018 368 047 523 842 11.01836804752384211.018\,368\,047\,523\,842\,11.018 368 047 523 842 1 67.109 721 899 032 5967.1097218990325967.109\,721\,899\,032\,5967.109 721 899 032 59 44.712 835 550 308 2344.7128355503082344.712\,835\,550\,308\,2344.712 835 550 308 23
$\bm{W}_{\text{in}}$ LoRA 0.681 224 489 192 509 40.68122448919250940.681\,224\,489\,192\,509\,40.681 224 489 192 509 4 67.064 607 143 402 167.064607143402167.064\,607\,143\,402\,167.064 607 143 402 1 43.031 200 766 563 41643.03120076656341643.031\,200\,766\,563\,41643.031 200 766 563 416
$\bm{W}_{\text{out}}$ LoRA 0.341 776 376 784 720 950.341776376784720950.341\,776\,376\,784\,720\,950.341 776 376 784 720 95 66.796 505 451 202 3966.7965054512023966.796\,505\,451\,202\,3966.796 505 451 202 39 42.278 078 198 432 9242.2780781984329242.278\,078\,198\,432\,9242.278 078 198 432 92
S6 All Full 4.169 116 475 966 0684.1691164759660684.169\,116\,475\,966\,0684.169 116 475 966 068 65.723 252 296 447 7565.7232522964477565.723\,252\,296\,447\,7565.723 252 296 447 75 39.696 556 329 727 1739.6965563297271739.696\,556\,329\,727\,1739.696 556 329 727 17
LoRA 0.383 152 556 698 126 50.38315255669812650.383\,152\,556\,698\,126\,50.383 152 556 698 126 5 64.183 932 542 800 964.183932542800964.183\,932\,542\,800\,964.183 932 542 800 9 40.119 516 849 517 8240.1195168495178240.119\,516\,849\,517\,8240.119 516 849 517 82
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},\bm{W}_{\bm{\Delta}}$ Full 4.001 065 760 075 972 54.00106576007597254.001\,065\,760\,075\,972\,54.001 065 760 075 972 5 65.965 962 409 973 1465.9659624099731465.965\,962\,409\,973\,1465.965 962 409 973 14 36.188 501 119 613 6536.1885011196136536.188\,501\,119\,613\,6536.188 501 119 613 65
LoRA 0.383 152 556 698 126 50.38315255669812650.383\,152\,556\,698\,126\,50.383 152 556 698 126 5 64.831 387 996 673 5864.8313879966735864.831\,387\,996\,673\,5864.831 387 996 673 58 39.472 505 450 248 7239.4725054502487239.472\,505\,450\,248\,7239.472 505 450 248 72
Table 11: Full benchmark results of LoRA on DART (Nan et al., 2021) dataset using Mamba-II 130M.
Layer Method # Params (%) All Easy Medium Hard Extra
All All Full $100.0$ 64.796 906 709 671 0264.7969067096710264.796\,906\,709\,671\,0264.796 906 709 671 02 85.887 098 312 377 9385.8870983123779385.887\,098\,312\,377\,9385.887 098 312 377 93 65.695 065 259 933 4765.6950652599334765.695\,065\,259\,933\,4765.695 065 259 933 47 54.022 985 696 792 654.022985696792654.022\,985\,696\,792\,654.022 985 696 792 6 42.168 673 872 947 6942.1686738729476942.168\,673\,872\,947\,6942.168 673 872 947 69
LoRA 0.706 416 999 407 535 90.70641699940753590.706\,416\,999\,407\,535\,90.706 416 999 407 535 9 64.506 769 180 297 8564.5067691802978564.506\,769\,180\,297\,8564.506 769 180 297 85 81.048 387 289 047 2481.0483872890472481.048\,387\,289\,047\,2481.048 387 289 047 24 66.367 715 597 152 7166.3677155971527166.367\,715\,597\,152\,7166.367 715 597 152 71 56.896 549 463 272 09556.89654946327209556.896\,549\,463\,272\,09556.896 549 463 272 095 42.771 083 116 531 3742.7710831165313742.771\,083\,116\,531\,3742.771 083 116 531 37
Linear Projection Matrices $\bm{W}_{\text{in}},\bm{W}_{\text{out}}$ LoRA 0.523 963 841 037 011 80.52396384103701180.523\,963\,841\,037\,011\,80.523 963 841 037 011 8 50.386 846 065 521 2450.3868460655212450.386\,846\,065\,521\,2450.386 846 065 521 24 68.548 387 289 047 2468.5483872890472468.548\,387\,289\,047\,2468.548 387 289 047 24 52.017 939 090 728 7652.0179390907287652.017\,939\,090\,728\,7652.017 939 090 728 76 44.827 586 412 429 8144.8275864124298144.827\,586\,412\,429\,8144.827 586 412 429 81 24.698 795 378 208 1624.6987953782081624.698\,795\,378\,208\,1624.698 795 378 208 16
$\bm{W}_{\text{in}}$ LoRA 0.349 920 379 444 911 50.34992037944491150.349\,920\,379\,444\,911\,50.349 920 379 444 911 5 57.543 522 119 522 09557.54352211952209557.543\,522\,119\,522\,09557.543 522 119 522 095 76.209 676 265 716 5576.2096762657165576.209\,676\,265\,716\,5576.209 676 265 716 55 59.417 039 155 960 0859.4170391559600859.417\,039\,155\,960\,0859.417 039 155 960 08 48.850 575 089 454 6548.8505750894546548.850\,575\,089\,454\,6548.850 575 089 454 65 33.734 938 502 311 7133.7349385023117133.734\,938\,502\,311\,7133.734 938 502 311 71
$\bm{W}_{\text{out}}$ LoRA 0.175 266 836 912 838 980.175266836912838980.175\,266\,836\,912\,838\,980.175 266 836 912 838 98 57.930 368 185 043 33557.93036818504333557.930\,368\,185\,043\,33557.930 368 185 043 335 81.048 387 289 047 2481.0483872890472481.048\,387\,289\,047\,2481.048 387 289 047 24 56.726 455 688 476 5656.7264556884765656.726\,455\,688\,476\,5656.726 455 688 476 56 51.724 135 875 701 90451.72413587570190451.724\,135\,875\,701\,90451.724 135 875 701 904 33.132 529 258 728 0333.1325292587280333.132\,529\,258\,728\,0333.132 529 258 728 03
S6 All Full 2.419 408 304 585 2852.4194083045852852.419\,408\,304\,585\,2852.419 408 304 585 285 55.125 725 269 317 6355.1257252693176355.125\,725\,269\,317\,6355.125 725 269 317 63 76.209 676 265 716 5576.2096762657165576.209\,676\,265\,716\,5576.209 676 265 716 55 56.053 811 311 721 856.053811311721856.053\,811\,311\,721\,856.053 811 311 721 8 42.528 736 591 339 1142.5287365913391142.528\,736\,591\,339\,1142.528 736 591 339 11 34.337 350 726 127 62534.33735072612762534.337\,350\,726\,127\,62534.337 350 726 127 625
LoRA 0.184 378 487 046 726 70.18437848704672670.184\,378\,487\,046\,726\,70.184 378 487 046 726 7 54.061 895 608 901 9854.0618956089019854.061\,895\,608\,901\,9854.061 895 608 901 98 74.193 549 156 188 9674.1935491561889674.193\,549\,156\,188\,9674.193 549 156 188 96 58.071 750 402 450 5658.0717504024505658.071\,750\,402\,450\,5658.071 750 402 450 56 45.977 011 322 975 1645.9770113229751645.977\,011\,322\,975\,1645.977 011 322 975 16 21.686 747 670 173 64521.68674767017364521.686\,747\,670\,173\,64521.686 747 670 173 645
${\bm{A}_{\text{log}}}$ Full 0.000 228 612 709 494 971 640.000228612709494971640.000\,228\,612\,709\,494\,971\,640.000 228 612 709 494 971 64 21.470 019 221 305 84721.47001922130584721.470\,019\,221\,305\,84721.470 019 221 305 847 45.967 742 800 712 58545.96774280071258545.967\,742\,800\,712\,58545.967 742 800 712 585 18.834 081 292 152 40518.83408129215240518.834\,081\,292\,152\,40518.834 081 292 152 405 11.494 252 830 743 7911.4942528307437911.494\,252\,830\,743\,7911.494 252 830 743 79 2.409 638 464 450 8362.4096384644508362.409\,638\,464\,450\,8362.409 638 464 450 836
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},\bm{W}_{\bm{\Delta}}$ Full 2.340 994 145 228 512.340994145228512.340\,994\,145\,228\,512.340 994 145 228 51 50.290 137 529 373 1750.2901375293731750.290\,137\,529\,373\,1750.290 137 529 373 17 72.983 872 890 472 4172.9838728904724172.983\,872\,890\,472\,4172.983 872 890 472 41 52.242 153 882 980 3552.2421538829803552.242\,153\,882\,980\,3552.242 153 882 980 35 39.655 172 824 859 6239.6551728248596239.655\,172\,824\,859\,6239.655 172 824 859 62 22.289 156 913 757 32422.28915691375732422.289\,156\,913\,757\,32422.289 156 913 757 324
LoRA 0.184 378 487 046 726 70.18437848704672670.184\,378\,487\,046\,726\,70.184 378 487 046 726 7 55.512 571 334 838 8755.5125713348388755.512\,571\,334\,838\,8755.512 571 334 838 87 77.419 352 531 433 177.419352531433177.419\,352\,531\,433\,177.419 352 531 433 1 55.156 952 142 715 45455.15695214271545455.156\,952\,142\,715\,45455.156 952 142 715 454 46.551 725 268 363 9546.5517252683639546.551\,725\,268\,363\,9546.551 725 268 363 95 33.132 529 258 728 0333.1325292587280333.132\,529\,258\,728\,0333.132 529 258 728 03
Table 12: Full benchmark results on the Spider (Yu et al., 2018) dataset using Mamba-II 1.3B.
Layer Method # Params (%) METEOR BLEU
All All Full $100.0$ $70.79$ $45.04$
Attention All LoRA 0.016 7060.0167060.016\,7060.016 706 $63.47$ $19.67$
MLP All LoRA 1.369 0841.3690841.369\,0841.369 084 $70.87$ $46.2$
Linear Projection Matrices + S6 All LoRA 0.308 3140.3083140.308\,3140.308 314 $70.16$ $39.99$
Linear Projection Matrices $\bm{W}_{\text{in}}$ LoRA 0.107 8460.1078460.107\,8460.107 846 $68.85$ $37.76$
$\bm{W}_{\text{out}}$ LoRA 0.053 9520.0539520.053\,9520.053 952 $67.67$ $31.85$
S6 All Full 0.535 3140.5353140.535\,3140.535 314 $69.23$ $35.49$
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ LoRA 0.147 1070.1471070.147\,1070.147 107 $66.55$ $24.16$
Table 13: Full benchmark results on DART (Nan et al., 2021) dataset using Jamba-Tiny-319M.
C.3 Limitations of Applying Input-injection Methods on SSMs
Task RTE MRPC CoLA SST-2 QNLI QQP MNLI Avg. Score
Prompt Tuning 56.0 71.6 12.0 89.4 76.8 79.6 61.5 63.8
Prefix-Tuning 69.5 75.7 43.4 91.5 83.4 83.1 35.6 68.6
Initial State Tuning 66.8 75.1 52.4 92.4 86.4 86.1 78.5 76.8
LoRA (Linear Projection Matrices) 70.4 82.8 60.6 92.4 88.4 87.7 81.5 80.5
Table 14: Comparison of prompt-tuning, prefix-tuning, initial state tuning, and LoRA on seven tasks from the GLUE benchmark. We report Matthews correlation ($\uparrow$) for CoLA, overall (matched and mismatched) accuracy ()($\uparrow$)( ↑ ) for MNLI, and accuracy for other tasks. Initial state tuning and LoRA are constrained to use less than 0.5% trainable parameters. Bold numbers indicate the best performance across all three methods, while underlined numbers show the highest score among input-injection methods (prefix-tuning and initial state tuning). Initial state tuning outperforms prefix-tuning and prompt-tuning on five out of seven tasks, while LoRA consistently outperforms all input-injection methods.